OpenStudy (anonymous):

solve the eqaution, 2sin^2(theta) = sin(theta)+1

OpenStudy (anonymous):

bring everything to one side and treat it as a quadratic equation: 2x^2 - x - 1 = 0

sam (.sam.):

bring over then factor, 2sin^2(theta) = sin(theta)+1 2sin^2(theta)- sin(theta)-1= 0 (sinx-1)(2sinx+1)=0 sinx=1 , sinx=-1/2

OpenStudy (anonymous):

Thanks Alot